# Efficient Reasoning Models: A Survey

### v1

@@ Abstract
Reasoning models have demonstrated remarkable progress in solving complex and logic-intensive tasks by generating extended Chain-of-Thoughts (CoTs) prior to arriving at a final answer. Yet, the emergence of this “slow-thinking” paradigm, with numerous tokens generated in sequence, inevitably introduces substantial computational overhead. To this end, it highlights an urgent need for effective acceleration. This survey aims to provide a comprehensive overview of recent advances in efficient reasoning. It categorizes existing works into three key directions: (1) shorter – compressing lengthy CoTs into concise yet effective reasoning chains; (2) smaller – developing compact language models with strong reasoning capabilities through techniques such as knowledge distillation, other model compression techniques, and reinforcement learning; and (3) faster – designing efficient decoding strategies to accelerate inference. A curated collection of papers discussed in this survey is available in our GitHub repository [^1] https://github.com/fscdc/Awesome-Efficient-Reasoning-Models. We will keep it updated with new research..
@@ 1 Introduction
Figure 1: Overview of efficient reasoning. We categorize existing efficient reasoning methods into three key directions based on how they improve reasoning efficiency: (1) make long CoT short (shorter); (2) build small language model with strong reasoning ability (smaller); and (3) let decoding more efficient (faster).
@@
Recent reasoning-oriented models, or Large Reasoning Models (LRMs) (Guo et al., 2025; Jaech et al., 2024), have achieved remarkable performance on complex reasoning tasks by generating long Chain-of-Thoughts (CoTs), enabling effective problem-solving in domains such as mathematics and coding (Sprague et al., 2024). However, while LRMs significantly improve performance on reasoning tasks, they also cause substantial overhead. Compared to standard LLMs, reasoning models lead to redundancy across multiple dimensions.
@@
A salient characteristic of reasoning models is their tendency to overthink by generating excessively long reasoning chains (Chen et al., 2024c; Sui et al., 2025a), which has naturally motivated efforts to improve efficiency by shortening reasoning paths. Meanwhile, recent studies (Wu et al., 2025c; Yang et al., 2025c; Jin et al., 2024b) challenge the assumption that longer CoTs always lead to better performance, showing even negative returns. To address this kind of CoT length redundancy, a range of methods have been proposed: reinforcement learning (RL) with length penalties (Luo et al., 2025a; Aggarwal & Welleck, 2025), supervised fine-tuning (SFT) on variable-length CoT data (Ma et al., 2025; Xia et al., 2025), and prompt-driven strategies that either guide reasoning paths or route inputs to more efficient solutions (Ding et al., 2024; Aytes et al., 2025). Furthermore, latent reasoning performs the process in latent space without generating explicit CoTs, making reasoning chains more concise (Hao et al., 2024; Su et al., 2025).
@@
In addition to excessively long reasoning chains, reasoning models typically rely on large model sizes to achieve strong reasoning performance (e.g., DeepSeek R1 (Guo et al., 2025) has 685B parameters), which leads to substantial computational and memory costs. To address this, model compression (Han et al., 2016) has proven effective in reducing model size redundancy in standard LLMs, naturally inspiring interest in how these techniques (e.g., distillation (Hinton et al., 2015), quantization (Gray & Neuhoff, 1998), and pruning (LeCun et al., 1989)) can be applied to improve reasoning efficiency. In parallel, another line of works directly builds small language models with strong reasoning abilities using RL (Li et al., 2023a; 2025d; Zhu et al., 2024b).
@@
Beyond length and model size redundancy, inefficiency can also arise during the decoding stage. To tackle this issue, a growing body of work focuses on accelerating inference through more efficient decoding strategies. Test-time scaling (TTS) strategies, while enhancing reasoning performance (Snell et al., 2024), also introduce latency redundancy during the decoding stage. Some methods (Sun et al., 2024a; Wang et al., 2024b) specifically target and optimize the speed of certain TTS strategies (Wang et al., 2022). Other approaches, like parallel decoding (Ning et al., 2023) and problem decomposition (Teng et al., 2025), also mitigate inefficiency.
@@
This survey aims to provide an overview of research in efficient reasoning. As illustrated in Figure 1, we categorize existing works into three key directions based on the type of redundancy they target: (1) making long CoT short (shorter), which focuses on enabling models to produce shorter reasoning paths while maintaining performance; (2) building small language model with strong reasoning abilities (smaller), which aims to develop compact language models with strong reasoning abilities; (3) making decoding more efficient (faster), which explores strategies to reduce latency during the decoding stage.
@@
The following sections of this survey cover the content as outlined below. Section 2 will explore key backgrounds closely related to efficient reasoning. Section 3 will systematically introduce various methods and their relationships across three categories: making long CoT short (see Section 3.1), building small language model with strong reasoning abilities (see Section 3.2), and letting decoding more efficient (see Section 3.3). Section 4 presents the evaluation metrics, as well as datasets and benchmarks. Section 5 will discuss the key challenges in the field and propose some potential future research directions, while Section 6 will conclude the survey. Additionally, Figure 2 illustrates the taxonomy of efficient reasoning methods discussed in this survey.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=, where level=2text width=6.3em,font=, where level=3text width=6.6em,font=, where level=4text width=6.6em,font=, [ Efficient Reasoning, ver, fill=blue!10, edge=blue!40 [ Make Long CoT
Short (Shorter)
(§$\S$3.1) , fill=cyan!20, edge=cyan!60 [ RL-based
(§$\S$3.1.1) , fill=cyan!20, edge=cyan!60 [ L1 (Aggarwal & Welleck, 2025), O1-Pruner (Luo et al., 2025a), DAST (Shen et al., 2025b),
THINKPRUNE (Hou et al., 2025), Kimi k1.5 (Team et al., 2025),
AGPO (Li et al., 2025a), Think When You Need (Yang et al., 2025b) , leaf, text width=41.2em, fill=cyan!20, edge=cyan!60 ] ] [ SFT-based
(§$\S$3.1.2) , fill=cyan!20, edge=cyan!60 [ TokenSkip (Xia et al., 2025), Distill2-to-1 (Yu et al., 2024), C3oT (Kang et al., 2024),
SPIRIT (Cui et al., 2025), TALE (Han et al., 2024), TOPS (Yang et al., 2025c),
CoT-Valve (Ma et al., 2025), LLM-Skip (Liu et al., 2024b), Self-training (Munkhbat et al., 2025) , leaf, text width=41.2em, fill=cyan!20, edge=cyan!60 ] ] [ Prompt-driven
(§$\S$3.1.3) , fill=cyan!20, edge=cyan!60 [ Prompt-guided , fill=cyan!20, edge=cyan!60 [ Concise CoT (Renze & Guven, 2024), MARP (Chen et al., 2024a), CoD (Xu et al., 2025b),
Break the Chain (Ding et al., 2024), Token-Complexity (Lee et al., 2025) , leaf, text width=33em, fill=cyan!20, edge=cyan!60 ] ] [ Prompt
attribute-aware , fill=cyan!20, edge=cyan!60 [ Claude 3.7 Sonnet (Anthropic., 2025), RouteLLM (Ong et al., 2024),
Sketch-of-Thought (Aytes et al., 2025), Self-REF (Chuang et al., 2024),
Confident or Seek Stronger (Chuang et al., 2025) , leaf, text width=33em, fill=cyan!20, edge=cyan!60 ] ] ] [ Latent Reasoning
(§$\S$3.1.4) , fill=cyan!20, edge=cyan!60 [ Explicit CoT to
Implicit CoT , fill=cyan!20, edge=cyan!60 [ Implicit-KD (Deng et al., 2023), Distill2-to-1 (Yu et al., 2024), CODI (Shen et al., 2025c),
LightThinker (Zhang et al., 2025a), CCoT (Cheng & Van Durme, 2024),
HCoT (Liu et al., 2024c), SoftCoT (Xu et al., 2025c), SI (Deng et al., 2024),
RELAY (Yu et al., 2025a), Reasoning with Latent Thoughts (Saunshi et al., 2025) , leaf, text width=33em, fill=cyan!20, edge=cyan!60 ] ] [ Latent Space
Representations , fill=cyan!20, edge=cyan!60 [ Implicit-KD (Deng et al., 2023), Distill2-to-1 (Yu et al., 2024), CODI (Shen et al., 2025c),
LightThinker (Zhang et al., 2025a), CCoT (Cheng & Van Durme, 2024),
HCoT (Liu et al., 2024c), SoftCoT (Xu et al., 2025c), SI (Deng et al., 2024),
RELAY (Yu et al., 2025a), Reasoning with Latent Thoughts (Saunshi et al., 2025),
Planning-Token (Wang et al., 2024c), Filler-Token (Pfau et al., 2024),
Coconut (Hao et al., 2024), Heima (Shen et al., 2025a),
Token Assorted (Su et al., 2025), Disentangling-Memory-and-Reasoning (Jin et al., 2024a) , leaf, text width=33em, fill=cyan!20, edge=cyan!60 ] ] ] ] [ Build Small
Language Model
with Strong
Reasoning Ability
(Smaller) (§$\S$3.2) , fill=green!20, edge=green!60 [ Distillation-based
(§$\S$3.2.1) , fill=green!20, edge=green!60 [ CoT-KD (Magister et al., 2022), MD (Li et al., 2023a), CD (Feng et al., 2024), Mix (Li et al., 2025d),
NAT (Li et al., 2024a), SCORE (Zhang et al., 2024), FDD (Zhu et al., 2024b),
SKIntern (Liao et al., 2025b), PRR (Zhao et al., 2024), ATM (Chen et al., 2024b),
Thinking slow, fast (Paliotta et al., 2025), Distill2-to-1 (Yu et al., 2024), DLCoT (Luo et al., 2025c) , leaf, text width=41.2em, fill=green!20, edge=green!60 ] ] [ Quantization &
Pruning (§$\S$3.2.2) , fill=green!20, edge=green!60 [ SLM-Foresee (Srivastava et al., 2025), Quantization Hurts Reasoning? (Liu et al., 2025c),
When Reasoning Meets Compression Zhang et al. (2025b) , leaf, text width=41.2em, fill=green!20, edge=green!60 ] ] [ RL-based
(§$\S$3.2.3) , fill=green!20, edge=green!60 [ Open-RS (Dang & Ngo, 2025), SimpleRL-Zoo (Zeng et al., 2025), DeepScaleR (Luo et al., 2025b) , leaf, text width=41.2em, fill=green!20, edge=green!60 ] ] ] [ Let Decoding
More Efficient
(Faster) (§$\S$3.3) , fill=purple!20, edge=purple!60 [ Efficiency for
TTS strategy
(§$\S$3.3.1) , fill=purple!20, edge=purple!60 [ Efficient Sampling , fill=purple!20, edge=purple!60 [ Fast Best-of-N (Sun et al., 2024a), ST-BoN (Wang et al., 2025b), FastMCTS (Li et al., 2025b),
Predictive-Decoding (Ma et al., 2024), $\phi$-Decoding (Xu et al., 2025a),
DPTS (Ding et al., 2025), FETCH (Wang et al., 2025a), SGD (Jin et al., 2024c) , leaf, text width=33em, fill=purple!20, edge=purple!60 ] ] [ Efficient
Self-consistency , fill=purple!20, edge=purple!60 [ ASC (Aggarwal et al., 2023), ESC (Li et al., 2024b), DSC (Wang et al., 2024b),
RASC (Wan et al., 2024), RPC (Zhou et al., 2025), CISC (Taubenfeld et al., 2025),
Self-Calibration (Huang et al., 2025), Path-Consistency (Zhu et al., 2024a) , leaf, text width=33em, fill=purple!20, edge=purple!60 ] ] ] [ Other Methods
(§$\S$3.3.2) , fill=purple!20, edge=purple!60 [ AoT (Teng et al., 2025), DISC (Light et al., 2025), Meta-Reasoner (Sui et al., 2025b), AR (Liu et al., 2025b),
MRT (Qu et al., 2025), TTC-Optimal Scaling (Snell et al., 2024), RSD (Liao et al., 2025a),
SpecReason Pan et al. (2025) , leaf, text width=41.2em, fill=purple!20, edge=purple!60 ] ] ] ]
@@ 2.1 Chain-of-Thought Reasoning
CoT (Wei et al., 2022) serves as a baseline reasoning approach, enabling LLMs to generate a sequence of intermediate steps before reaching the final answer, thus significantly improving performance on complex reasoning tasks. Various extensions have subsequently been proposed to further enhance reasoning capabilities. For instance, Tree-of-Thought (ToT) (Yao et al., 2023) generalizes the linear CoT structure into a tree, facilitating the exploration of multiple reasoning paths through backtracking and lookahead strategies. Graph-of-Thoughts (GoT) (Besta et al., 2024) has expanded this approach into graph structures to better capture dependencies and compositional relationships among reasoning steps, substantially improving reasoning quality. Additionally, some specialized CoT variants are task-specific. PoT (Chen et al., 2022) disentangles reasoning from computation by having the language model generate programmatic reasoning steps (i.e., expressing thoughts as code), which an external calculator executes to obtain the final answer, making this approach particularly effective for math and financial reasoning tasks. CoS (Hu et al., 2024), on the other hand, targets spatial reasoning by leveraging compressed symbolic representations of spatial relations to reduce token usage.
Scaling test-time computation (TTC) is another road for enhancing reasoning performance. For instance, Best-of-N selects the top-scoring answer from multiple samples, whereas self-consistency chooses the most consistent answer across multiple reasoning chains. Initially, the focus of TTS strategies was primarily on maximizing performance, often at the cost of efficiency. As the demand grew, more recent approaches tried to find ways to improve efficiency without significantly sacrificing performance. Fast Best-of-N (Sun et al., 2024a) introduces speculative rejection during sampling to proactively discard low-quality reasoning paths. SoT (Ning et al., 2023) employs a two-stage decoding strategy by generating a skeleton and filling nodes in parallel. Additionally, an empirical study (Wu et al., 2025b) investigates the trade-offs between the efficiency and performance of various TTS strategies (e.g., Best-of-N, weighted voting) under different model sizes and computation budgets, providing practical insights for further research and deployment.

### v2

@@ Abstract
Reasoning models have demonstrated remarkable progress in solving complex and logic-intensive tasks by generating extended Chain-of-Thoughts (CoTs) prior to arriving at a final answer. Yet, the emergence of this “slow-thinking” paradigm, with numerous tokens generated in sequence, inevitably introduces substantial computational overhead. To this end, it highlights an urgent need for effective acceleration. This survey aims to provide a comprehensive overview of recent advances in efficient reasoning. It categorizes existing works into three key directions: (1) shorter – compressing lengthy CoTs into concise yet effective reasoning chains; (2) smaller – developing compact language models with strong reasoning capabilities through techniques such as knowledge distillation, other model compression techniques, and reinforcement learning; and (3) faster – designing efficient decoding strategies to accelerate inference of reasoning models. A curated collection of papers discussed in this survey is available in our GitHub repository: https://github.com/fscdc/Awesome-Efficient-Reasoning-Models.
@@ 1 Introduction
Recent reasoning-oriented models, or Large Reasoning Models (LRMs) (Guo et al., 2025; Jaech et al., 2024), have achieved remarkable performance on complex reasoning tasks by generating long Chain-of-Thoughts (CoTs), enabling effective problem-solving in domains such as mathematics and coding (Sprague et al., 2024). However, while LRMs significantly improve performance on reasoning tasks, they also cause substantial overhead. Compared to standard Large Language Models (LLMs), reasoning models lead to redundancy across multiple dimensions.
Figure 1: Overview of efficient reasoning. We categorize existing efficient reasoning methods into three key directions based on how they improve reasoning efficiency: (1) make long CoT short (shorter); (2) build small language models with strong reasoning ability (smaller); and (3) let decoding more efficient (faster).
@@
A salient characteristic of reasoning models is their tendency to overthink by generating excessively long reasoning chains (Chen et al., 2024c; Sui et al., 2025a), which has naturally motivated efforts to improve efficiency by shortening reasoning paths. Meanwhile, recent studies (Wu et al., 2025d; Yang et al., 2025c; Jin et al., 2024b) challenge the assumption that longer CoTs always lead to better performance, showing even negative returns. To address this kind of CoT length redundancy, a range of methods have been proposed: reinforcement learning (RL) with length penalty (Luo et al., 2025a; Aggarwal & Welleck, 2025), supervised fine-tuning (SFT) on variable-length CoT data (Ma et al., 2025; Xia et al., 2025), and prompt-driven strategies that either guide reasoning paths or route inputs to more efficient solutions (Ding et al., 2024; Aytes et al., 2025). Furthermore, latent reasoning performs the process in latent space without generating explicit CoTs, making reasoning chains more concise (Hao et al., 2024; Su et al., 2025).
@@
In addition to excessively long reasoning chains, reasoning models typically rely on large model sizes to achieve strong reasoning performance (e.g., DeepSeek R1 (Guo et al., 2025) has 685B parameters), which leads to substantial computational and memory costs. To address this, model compression (Han et al., 2016) has proven effective in reducing model size redundancy in standard LLMs, naturally inspiring interest in how these techniques (e.g., distillation (Hinton et al., 2015), quantization (Gray & Neuhoff, 1998), and pruning (LeCun et al., 1989)) can be applied to improve reasoning efficiency. In parallel, another line of work directly builds small language models with strong reasoning abilities using RL (Li et al., 2023a; 2025e; Zhu et al., 2024b).
@@
Beyond length and model size redundancy, inefficiency can also arise during the decoding stage. A growing body of work focuses on accelerating inference through more efficient decoding strategies to tackle this issue. Test-time scaling (TTS) strategies, while enhancing reasoning performance (Snell et al., 2024), also introduce latency redundancy during the decoding stage. Some methods (Sun et al., 2024a; Wang et al., 2024b) specifically target and optimize the speed of certain TTS strategies (Wang et al., 2022a). Other approaches, like parallel decoding (Ning et al., 2023) and problem decomposition (Teng et al., 2025), also mitigate inefficiency.
@@
This survey aims to provide an overview of research in efficient reasoning. As illustrated in Figure 1, we categorize existing works into three key directions based on the type of redundancy they target: (1) making long CoT short (shorter), which focuses on enabling models to produce shorter reasoning paths while maintaining performance; (2) building small language model with strong reasoning abilities (smaller), which aims to endow compact models with the ability to solve complex reasoning tasks; (3) making decoding more efficient (faster), which explores strategies to reduce latency during the decoding stage.
@@
The following sections of this survey cover the content as outlined below. Section 2 will explore key backgrounds closely related to efficient reasoning. Section 3 will systematically introduce various methods and their relationships across three categories. Section 4 presents the evaluation metrics, as well as datasets and benchmarks. Section 5 will discuss the key challenges in the field and propose some potential future research directions, while Section 6 will conclude the survey. Additionally, Figure 2 illustrates the taxonomy of efficient reasoning methods discussed in this survey.
[…]
Efficient Reasoning
Make Long CoT Short (Shorter) ($\S$3.1) RL-based ($\S$3.1.1) L1 (Aggarwal & Welleck, 2025), O1-Pruner (Luo et al., 2025a), DAST (Shen et al., 2025b), THINKPRUNE (Hou et al., 2025), Kimi k1.5 (Team et al., 2025), AGPO (Li et al., 2025a), Think When You Need (Yang et al., 2025b) SFT-based ($\S$3.1.2) TokenSkip (Xia et al., 2025), Distill2-to-1 (Yu et al., 2024), C3oT (Kang et al., 2024), SPIRIT (Cui et al., 2025), TALE (Han et al., 2024), TOPS (Yang et al., 2025c), CoT-Valve (Ma et al., 2025), LLM-Skip (Liu et al., 2024b), Self-training (Munkhbat et al., 2025) Prompt-driven ($\S$3.1.3) Prompt-guided Concise CoT (Renze & Guven, 2024), MARP (Chen et al., 2024a), CoD (Xu et al., 2025c), Break the Chain (Ding et al., 2024), Token-Complexity (Lee et al., 2025) Prompt attribute-aware Claude 3.7 Sonnet (Anthropic., 2025), RouteLLM (Ong et al., 2024), Sketch-of-Thought (Aytes et al., 2025), Self-REF (Chuang et al., 2024), Confident or Seek Stronger (Chuang et al., 2025) Latent Reasoning ($\S$3.1.4) Explicit CoT to Implicit CoT Implicit-KD (Deng et al., 2023), Distill2-to-1 (Yu et al., 2024), CODI (Shen et al., 2025c), LightThinker (Zhang et al., 2025a), CCoT (Cheng & Van Durme, 2024), HCoT (Liu et al., 2024c), SoftCoT (Xu et al., 2025d), SI (Deng et al., 2024), RELAY (Yu et al., 2025a), Reasoning with Latent Thoughts (Saunshi et al., 2025) Latent Space Representations Implicit-KD (Deng et al., 2023), Distill2-to-1 (Yu et al., 2024), CODI (Shen et al., 2025c), LightThinker (Zhang et al., 2025a), CCoT (Cheng & Van Durme, 2024), HCoT (Liu et al., 2024c), SoftCoT (Xu et al., 2025d), SI (Deng et al., 2024), RELAY (Yu et al., 2025a), Reasoning with Latent Thoughts (Saunshi et al., 2025), Planning-Token (Wang et al., 2024c), Filler-Token (Pfau et al., 2024), Coconut (Hao et al., 2024), Heima (Shen et al., 2025a), Token Assorted (Su et al., 2025), Disentangling-Memory-and-Reasoning (Jin et al., 2024a) Build Small Language Model with Strong Reasoning Ability (Smaller) ($\S$3.2) Distillation-based ($\S$3.2.1) CoT-KD (Magister et al., 2022), MD (Li et al., 2023a), CD (Feng et al., 2024c), Mix (Li et al., 2025e), NAT (Li et al., 2024a), SCORE (Zhang et al., 2024), FDD (Zhu et al., 2024b), SKIntern (Liao et al., 2025b), PRR (Zhao et al., 2024), ATM (Chen et al., 2024b), Thinking slow, fast (Paliotta et al., 2025), Distill2-to-1 (Yu et al., 2024), DLCoT (Luo et al., 2025c), M1 (Wang et al., 2025b), Quantization & Pruning ($\S$3.2.2) SLM-Foresee (Srivastava et al., 2025), Quantization Hurts Reasoning? (Liu et al., 2025c), When Reasoning Meets Compression (Zhang et al., 2025b) RL-based ($\S$3.2.3) Open-RS (Dang & Ngo, 2025), SimpleRL-Zoo (Zeng et al., 2025a), DeepScaleR (Luo et al., 2025b) Let Decoding More Efficient (Faster) ($\S$3.3) Efficiency for TTS strategy ($\S$3.3.1) Efficient Sampling Fast Best-of-N (Sun et al., 2024a), ST-BoN (Wang et al., 2025d), FastMCTS (Li et al., 2025b), Predictive-Decoding (Ma et al., 2024), $\phi$-Decoding (Xu et al., 2025a), DPTS (Ding et al., 2025), FETCH (Wang et al., 2025a) Efficient Self-consistency ASC (Aggarwal et al., 2023), ESC (Li et al., 2024b), DSC (Wang et al., 2024b), RASC (Wan et al., 2024), RPC (Zhou et al., 2025), CISC (Taubenfeld et al., 2025), Self-Calibration (Huang et al., 2025), Path-Consistency (Zhu et al., 2024a) Other Methods ($\S$3.3.2) AoT (Teng et al., 2025), DISC (Light et al., 2025), Meta-Reasoner (Sui et al., 2025c), AR (Liu et al., 2025b), MRT (Qu et al., 2025b), TTC-Optimal Scaling (Snell et al., 2024), RSD (Liao et al., 2025a), SpecReason (Pan et al., 2025), SoT (Ning et al., 2023), SGD (Jin et al., 2024c)
@@ 2.1 Chain-of-Thought Reasoning
CoT (Wei et al., 2022) serves as a baseline reasoning approach, enabling LLMs to generate a sequence of intermediate steps before reaching the final answer, thus significantly improving performance on complex reasoning tasks. Various extensions have subsequently been proposed to further enhance reasoning capabilities. For instance, Tree-of-Thought (ToT) (Yao et al., 2023) generalizes the linear CoT structure into a tree, facilitating the exploration of multiple reasoning paths through backtracking and lookahead strategies. Graph-of-Thoughts (GoT) (Besta et al., 2024) has expanded this approach into graph structures to better capture dependencies and compositional relationships among reasoning steps, substantially improving reasoning quality. Additionally, some specialized CoT variants are task-specific. PoT (Chen et al., 2022) disentangles reasoning from computation by having the language model generate programmatic reasoning steps (i.e., expressing thoughts as code), which an external calculator executes to obtain the final answer, making this approach particularly effective for math and financial tasks. CoS (Hu et al., 2024), on the other hand, targets spatial reasoning by leveraging compressed symbolic representations of spatial relations to reduce token usage.
2.2 Reasoning Models and Underlying Techniques
Recent reasoning models have moved beyond early prompting-based CoT techniques by internalizing step-by-step reasoning through SFT and RL. Building structured reasoning paradigms mentioned in Section 2.1, these models are trained to generate reasoning traces aligned with human-like logic. RL plays a crucial role by optimizing for reasoning quality using reward signals based on correctness, format alignment, and process supervision (Xu et al., 2025b; Ouyang et al., 2022; Zhou et al., 2023). Advanced models like OpenAI o1 (OpenAI, 2024) are believed to incorporate tree-search strategies (Coulom, 2006) and process reward models to guide the exploration of intermediate steps. Others, such as DeepSeek R1 (Guo et al., 2025), employ rule-based reward functions to reinforce correct reasoning steps.
2.3 Test-Time Scaling
Scaling test-time computation (TTC) is another road for enhancing reasoning performance (Snell et al., 2024; Zeng et al., 2025b). Scaling can be approached from two complementary dimensions: horizontal and vertical. The horizontal perspective involves generating multiple samples and selecting the best answer. Best-of-N (Cobbe et al., 2021; Sun et al., 2024a) selects the top-scoring response, while self-consistency (Wang et al., 2022a) identifies the most consistent answer across reasoning chains. The vertical perspective focuses on increasing the length of a single reasoning path. For example, Self-Refine (Madaan et al., 2023) iteratively improves an initial response via self-evaluation, while other works (Chen et al., 2024d; Gou et al., 2024) leverage external feedback to guide the refinement process. Additionally, an empirical study (Wu et al., 2025c) investigates the trade-offs between the efficiency and performance of various TTS strategies (e.g., Best-of-N, weighted voting) under different model sizes and computation budgets, providing practical insights for further research and deployment.
2.4 Model Compression
Model compression strategies are widely used to reduce the size and computational overhead of models (Han et al., 2016). Common approaches include quantization (Gray & Neuhoff, 1998; Frantar et al., 2023a; Lin et al., 2024; Xiao et al., 2023), which reduces model size by lowering the precision of model parameters. Pruning (LeCun et al., 1989; Ma et al., 2023; Fang et al., 2023; Wang et al., 2021) removes less significant or redundant model parameters to achieve sparsity, reducing model size and inference latency. Unlike the above techniques, knowledge distillation (Hinton et al., 2015; Wang et al., 2022b; Liu et al., 2019) achieves compression not by directly modifying the original model, but by transferring knowledge from a larger, well-trained teacher model to a smaller student model, allowing the student to replicate the teacher’s behavior while maintaining comparable performance (see details about model compression in Appendix A.1).